15) A test for equality of two variances has sample sizes n1 = 13 and n2 = 10. The degrees of freedom for the test are
A.

21.
B.

13 and 10.
C.

12 and 9.
D.

26.

Answer by stanbon(75887)

15) A test for equality of two variances has sample sizes n1 = 13 and n2 = 10. The degrees of freedom for the test are
---
df = n1+n2-2 = 13+10-2 = 21
----------------------------------
Cheers,
Stan H.
----------------------------------
A.
21.
B.
13 and 10.
C.
12 and 9.
D.
26.